Bleach now come in a 8.5% concentration. I want to make 500ml of a 10% bleach solution using the 8.5% concentration. Need ot know the amount of bleach needed. Thank you!
Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

Bleach now come in a 8.5% concentration. I want to make 500ml of a 10% bleach solution using the 8.5% concentration. Need ot know the amount of bleach needed.
-------------
You are going to have to add some bleach solution that is more
than 10%.
I'll assume you have a 100% solution available for the mixture.
--------------------
active + active = active
0.085x + 1*(500-x) = 0.10*500
-----
85x + 1000*500 - 1000x = 100*500
-----
-915x = -900*500
x = 491.8 ml (amt. of 8.5% solution needed)
500-x = 8.2 ml (amt. of 100% concentrate needed)
====================
Cheers,
Stan H.
====================

Answer by josgarithmetic(39617)

That is not right. You want 10% bleach, but bleach comes available as 8.5% bleach. You must mix the 8.5% bleach with a bleach solution HIGHER in concentration than 10%. Is a 12.5% bleach solution available?

Correct your problem description and question, and then I could help.